Let the plane P pass through the 3 points R = (1, 1, 6), S= (2,5,4) and T = (1,2,3).


Fill in the following:

  1. RS = (__,__,__)
  2. RT = (__,__,__)
  3. RS x RT = (__,__,__)




1
Expert's answer
2021-07-07T11:18:54-0400

R = (1, 1, 6)

S = (2, 5, 4)

T = (1, 2, 3)


To check if the points are collinear

"\begin{vmatrix}
 1 & 1& 6\\
 2& 5& 4\\1&2&3
\end{vmatrix}=-1"


"\begin{vmatrix}
 (x- 1 )&( y-1)& (z-6)\\
 (1-2)& (1-5)& (6-4)\\(1-1)&(1-2)&(6-3)
\end{vmatrix}=0"


"\begin{vmatrix}
 (x- 1 )&( y-1)& (z-6)\\
 -1& -4& 2\\0&-1&3
\end{vmatrix}= 0"

On solving,

"-14(x-1)-((-3)(y-1))+1(z-6)= 0"

"-14x+14+3y-3+z-6 =0"

"14x-3y-z = 5" is the required equation.


R = 1i + 1j + 6k

S = 2i + 5j + 4k

T = 1i + 2i + 3k


1. "\overrightarrow{RS}= (x_2 – x_1) \hat{i} + (y_2 – y_1) \hat{j} + (z_2 – z_1) \hat{k}"

"\overrightarrow{RS}= (2 – 1) \hat{i} + (5–1) \hat{j} + (4– 6) \hat{k} = \hat{i}+ 4\hat{j}-2 \hat{k}= (1,4,-2)"


2.

"\overrightarrow{RT}= (x_3– x_1) \hat{i} + (y_3– y_1) \hat{j} + (z_3 – z_1) \hat{k}"

"\overrightarrow{RT}= (1– 1) \hat{i} + (2–1) \hat{j} + (3– 6) \hat{k} = 0\hat{i}+ \hat{j}-3\hat{k}= (0,1,-3)"


3.

"\overrightarrow{RS} ×\overrightarrow{ RT }= (\hat{i}+ 4\hat{j}-2 \hat{k})×(0\hat{i}+ \hat{j}-3\hat{k}) ="

"\begin{vmatrix}
 i & j& k\\
 1& 4& -2\\0&1&-3
\end{vmatrix}="

"i(-12+2) -j(-3-0)+k(1-0)=-10\hat{i}+ 3\hat{j}+1\hat{k}"


"\therefore\overrightarrow{RS} ×\overrightarrow{ RT }= (-10 ,3,1)"